OpenStudy (anonymous):

(-6) to the 5th power multiplied by (-6) to the 7th power. I need the answer

OpenStudy (anonymous):

Since we have the same base (which is -6) -> (-6)^5*(-6)^7 = (-6)^12 = 6^12. This is 2176782336, but I can't imagine that number is that useful...

OpenStudy (anonymous):

Nope. i don't think that is the right answer, but thank you

OpenStudy (anonymous):

I got the same as her. 2176782336

OpenStudy (anonymous):

\[(-6)^{5}*(-6)^{7}=(-7776)*(-279936)=2176782336\] collegemom, what do you think is incorrect with the preceding result?